OpenStudy (anonymous):

Find the minimum value for the function: y = x^2 + 9x + 18

OpenStudy (anonymous):

do u knwo how to complete the square? or do u factor it?

OpenStudy (anonymous):

lets factor its easy

OpenStudy (anonymous):

its under my quadratic function lesson so idk

OpenStudy (anonymous):

what 2 numbers when you multiply them =18 and when u add them =9?

OpenStudy (anonymous):

is there a number for that?

OpenStudy (anonymous):

6*3=18 6+3=9

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so is the answer 9 ?

OpenStudy (anonymous):

yaaaa factoring i just taught this to my class today. hmmmm nooooo we r factoring. We r trying to figure out 2 numbers that when u mulyiply them give us 18 and when u add them gives us 9

OpenStudy (anonymous):

finding the zeros is a long and complicated way to find the minimum

OpenStudy (anonymous):

umm well we can complete the square

OpenStudy (anonymous):

but i dont think she knows that

OpenStudy (zarkon):

use the vertex formula

OpenStudy (anonymous):

minimum is at the vertex, and first coordinate of the vertex of \[y=ax^2+bx+c\] is always \(-\frac{b}{2a}\)

OpenStudy (anonymous):

when doing the vertx thing... is the second number the minimum?

OpenStudy (anonymous):

in your case it is \(-\frac{9}{2}\) replace \(x\) by \(-\frac{9}{2}\) to find the minimum

OpenStudy (anonymous):

minimum means minimum value of \(y\) so it is the second coordinate, not the first

OpenStudy (anonymous):

so yes, what you said, the second number is the minimum.

OpenStudy (anonymous):

ohhh i like this method. We never learnt this stuff. Like the textbook always made us complete the square. oh cooll

OpenStudy (anonymous):

can i just plug it into my graphing calculator to get the anwser ?

OpenStudy (anonymous):

completing the square is for suckers, unless of course you are solving a quadratic equation

OpenStudy (lgbasallote):

last i heard minimum value was in a calculus class o.O i never knew it was in algebra too

OpenStudy (lukecrayonz):

It's in Algebra 2.

OpenStudy (lukecrayonz):

You can use a calculator to solve it

OpenStudy (anonymous):

hahah its in math 11 here in canada

OpenStudy (anonymous):

\[(-\frac{9}{2})^2+9\times (\frac{-9}{2})+18\] then a calculator if you need one

OpenStudy (anonymous):

you turn math up to eleven in canada? mine only goes up to ten

OpenStudy (lukecrayonz):

Ti-84, input the equation into Y= then click 2nd Trace, and do minimum. It will say "Left bound?", go to the left side of the graph and press enter. Right bound? Obviously, right side. It will say "Guess?", which means put it to where you think the minimum y value is.

OpenStudy (lukecrayonz):

So for this, you get x=-4.5 and y=-2.25

OpenStudy (anonymous):

sooo many numbers ... ahhhhh!!! >_< what about TI-83 plus?

OpenStudy (anonymous):

hahaha grade 11 math lol

OpenStudy (lgbasallote):

so taking the derivative and equating to zero is another story?

OpenStudy (lukecrayonz):

It'll be exactly the same basically I think.

OpenStudy (lukecrayonz):

Yes lgb haha

OpenStudy (anonymous):

they give us this freshman yr. imma senior now tryna make it up.. sucks :/

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

using the derivative to find the minimum value of a quadratic. like using a glock to kill a wingspanroach